Grade 8Algebra

If tan A + cot A = 4; then find tan^4 A + cot^4 A= ??

Shiphrah
8 Years agoGrade 8

2 Answers

Arun
ApprovedApproved Tutor Answer8 Years ago
Dear student
 
tan(a) + cot(a) = 4 

[tan(a) + cot(a)]² = 16 

tan²(a) + 2tan(a).cot(a) + cot²(a) = 16 

tan²(a) + 2 + cot²(a) = 16 

tan²(a) + cot²(a) = 14 

[tan²(a) + cot²(a)]² = 196 

tan^4(a) + 2tan²(a).cot²(a) + cot^4(a) = 196 

tan^4(a) + 2 + cot^4(a) = 196 

tan^4(a) + cot^4(a) = 194 
 
 
Regards
Arun (askIITians forum expert)
Arun
8 Years ago
Dear student
 
tan(a) + cot(a) = 4 
 
[tan(a) + cot(a)]² = 16 
 
tan²(a) + 2tan(a).cot(a) + cot²(a) = 16 
 
tan²(a) + 2 + cot²(a) = 16 
 
tan²(a) + cot²(a) = 14 
 
[tan²(a) + cot²(a)]² = 196 
 
tan^4(a) + 2tan²(a).cot²(a) + cot^4(a) = 196 
 
tan^4(a) + 2 + cot^4(a) = 196 
 
tan^4(a) + cot^4(a) = 194 
 
 
Regards
Arun (askIITians forum expert)